# Enhancing Tool Retrieval with Iterative Feedback from Large Language Models

### v1

@@ Abstract
Tool learning aims to enhance and expand large language models’ (LLMs) capabilities with external tools, which has gained significant attention recently. Current methods have shown that LLMs can effectively handle a certain amount of tools through in-context learning or fine-tuning. However, in real-world scenarios, the number of tools is typically extensive and irregularly updated, emphasizing the necessity for a dedicated tool retrieval component. Tool retrieval is nontrivial due to the following challenges: 1) complex user instructions and tool descriptions; 2) misalignment between tool retrieval and tool usage models. To address the above issues, we propose to enhance tool retrieval with iterative feedback from the large language model. Specifically, we prompt the tool usage model, i.e., the LLM, to provide feedback for the tool retriever model in multi-round, which could progressively improve the tool retriever’s understanding of instructions and tools and reduce the gap between the two standalone components. We build a unified and comprehensive benchmark to evaluate tool retrieval models. The extensive experiments indicate that our proposed approach achieves advanced performance in both in-domain evaluation and out-of-domain evaluation.
@@
Qiancheng Xu1, Yongqi Li1†, Heming Xia1, Wenjie Li1 1Department of Computing, The Hong Kong Polytechnic University, China {qiancheng.xu, he-ming.xia}@connect.polyu.hk liyongqi0@gmail.comcswjli@comp.polyu.edu.hk
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable success in language-related tasks and are considered a potential pathway to achieving artificial general intelligence Zhao et al. (2023). However, despite their powerful capabilities, LLMs are still limited in many aspects, such as knowledge update and mathematical reasoning. A promising way to overcome these limitations is to empower LLMs with external tools, known as tool learning. Tool learning not only enhances LLMs’ performance on existing tasks but also allows them to tackle tasks that were previously beyond their reach. Besides, the ability to use tools is a crucial hallmark on the path to advanced intelligence.
@@
Despite the practicality and necessity, tool retrieval has been inadequately studied. Some approaches have adopted traditional document retrieval methods to retrieve tools for LLMs Li et al. (2023); Patil et al. (2023); Qin et al. (2023b). However, we argue that they overlook the unique challenges of tool retrieval for LLMs: 1) Complex user instructions and tool descriptions. As illustrated in Figure 2, compared with document retrieval, user instructions are usually ambiguous and complex, and the reputation rate between instructions and corresponding tool descriptions is much lower. Unfortunately, the retriever model is typically limited in its capacities because of the efficiency requirements, which makes tool retrieval more difficult and challenging. 2) Misalignment between tool retrieval and tool usage models. Previous approaches deploy the tool retriever separately from the downstream tool-usage model, which hinders the LLM from knowing which tools are really useful from the tool-usage perspective. Thus, it will result in a tool recognition gap between the tool retriever and tool usage model, degrading the tool-use performance further.
@@ 2.1 Tool Learning in LLMs
Tool learning aims to equip LLMs with external tools to enhance and expand their capabilities Schick et al. (2023); Tang et al. (2023); Yao et al. (2023); Qin et al. (2023a); Shen et al. (2023). Generally, existing tool learning methods could be categorized into in-context learning and fine-tuning approaches. The former approach encourages LLMs to use tools with descriptions, documentation, or demonstrations, while the latter one trains the parameters of LLMs using specially created tool-use datasets. However, no matter whether the in-context learning or fine-tuning approach encounters severe challenges in real-world scenarios, where the candidate tools are extensive and frequently updated. Therefore, it is crucial to equip LLMs with a tool retrieval component to select appropriate tools from a large-scale tool set. Recent works have proposed a stopgap measure through traditional document retrieval methods Li et al. (2023); Patil et al. (2023); Qin et al. (2023b). In this work, we aim to develop a specialized method for retrieving tools.
@@ 4.1 Overview
Recent studies have found that LLMs show a great capability in acting as a critic Zheng et al. (2023) and could provide comprehensive feedback to improve performance across a range of tasks Madaan et al. (2023); Asai et al. (2023). Inspired by those observations, we propose an innovative framework that leverages the LLM’s feedback to improve the tool retrieval process iteratively.
@@ 4.2 Feedback Generation
The feedback from the LLM is finalized made on the current user instruction $q^{t}$ Specifically, we prompt the LLM to generate refined instruction with enriched information in two dimensions: 1) more detailed and personalized content about those user’s intent which have not been solved by current tools, helping the retriever explore other relevant tools; (2) more scenario-specific tool-usage information about existing appropriate tools, helping the retriever give higher ranking priority to those tools. This step can be formulated as,
@@ 4.3 Iteration-Aware Feedback Training
We concatenate a special token “Iteration $t$” in front of the instruction, where $t$ is the instruction’s iteration step (e.g., “Iteration $t-1$” for $q^{t-1}$ and “Iteration $t$” for $q^{t}$). The final training loss is formulated as the sum of losses in each iteration as follows,
$\mathcal{L}_{feedback}=\sum_{t=1}^{T}\alpha^{t}\mathcal{L}(q^{t}),$ (5)
where $\alpha^{t}$ is a balancing factor. In this way, the LLM’s comprehensive knowledge of the user requirements could be injected into the retriever through those refined instructions. Besides, with the aid of iteration-aware tokens and joint-training manner, the retriever could maintain a balance between newly learned knowledge and previously acquired knowledge.
We also employ the hard negative sampling in training. Concretely, for each given instruction, we randomly sample an incorrect tool from the retrieved top-$K$ tool list. The high similarity scores of those tools indicate that they are prone to be mistaken as correct tools by the retriever. In feedback training, we utilize those tool-instruction pairs as hard negative samples.
Then, the loss function for each iteration could be calculated as,
@@ 5.1 Setup
Datasets and evaluation. To assess the tool retrieval performance of models, we conduct a tool retrieval benchmark, referred to as TR-bench, based on three datasets, including ToolBench Qin et al. (2023b), T-Eval Chen et al. (2023), and UltraTools Huang et al. (2024). To address real-world requirements, we conduct evaluations in both in-domain and out-of-domain settings. Specifically, the training set is from ToolBench, while the test set of ToolBench is employed for in-domain evaluation, and the test sets from T-Eval and UltraTools are used for out-of-domain evaluation. The statistics of TR-bench are summarized in Table 1.
Following previous work Qin et al. (2023b), we adopt the Normalized Discounted Cumulative Gain (NDCG) Järvelin and Kekäläinen (2002), an ideal ranking metric for tool retrieval since it could evaluate the relevance and quality of retrieved tool candidates according to their ranked orders. We report NDCG@$m$ ($m=1,3,5,10$), where $m$ refers to top-$m$ ranked search results for evaluation.
@@
Out-of-domain Evaluation Teval 553 50
@@
Table 1: Statistics of the TR-bench, which is conducted from ToolBench Qin et al. (2023b), T-Eval Chen et al. (2023), and UltraTools Huang et al. (2024).
@@
Table 3: Out-of-domain evaluation on TR-bench in terms of NDCG@$m$ under two scenarios, T-Eval Chen et al. (2023) and UltraTools Huang et al. (2024). % improve represents the relative improvement achieved by our method over the previously best tool retrieval method.
Implementation details. We employ Sentence-BERT Reimers and Gurevych (2019) to train our retriever model based on BERT-base Devlin et al. (2019). We set the learning rate to $2e{-5}$ with 500 warm-up steps. The batch size in training is set to $64$. We utilize ChatGPT (gpt-3.5-turbo-0125) [^2] https://openai.com/index/introducing-chatgpt-and-whisper-apis/. as the LLM for giving feedback. The number of tool candidates $K$, the balancing factor $\alpha$, and the iteration round $T$ are set to 10, 1, and 3, respectively. We have trained the model several times to confirm that the improvement is not a result of random chance and present the mid one. Our experiments were conducted on four NVIDIA A6000 GPUs with 48 GB of memory
@@ 5.3 Ablation Study
Table 5: Analysis on iteration round under the intra-category multi-tool instructions (I2) scenario. The efficiency is measured by the time consumption to complete one user instruction.
@@
Table 6: Analysis on different base models under the intra-category multi-tool instructions (I2) scenario.
@@
Table 7: Analysis on embedding sizes under the intra-category multi-tool instructions (I2) scenario.
@@ 6 Conclusion and Future Work
1) Undoubtedly, our iterative refinement will reduce the inference speed of the tool retrieval. We have evaluated the efficiency as the number of iterative rounds increases. Fortunately, we observed that just one additional round of refinement could yield significant improvements. Furthermore, the performance enhancement of the tool retrieval is crucial for the subsequent tool usage model. 2) Similar to document retrieval, the used datasets in our work also contain “false negative” samples. For instance, some tools may be capable of handling the user’s instruction but are not labeled as positive. This can disrupt the training and evaluation of tool retrieval and is a common limitation in many retrieval scenarios.

### v2

@@ Abstract
Tool learning aims to enhance and expand large language models’ (LLMs) capabilities with external tools, which has gained significant attention recently. Current methods have shown that LLMs can effectively handle a certain amount of tools through in-context learning or fine-tuning. However, in real-world scenarios, the number of tools is typically extensive and irregularly updated, emphasizing the necessity for a dedicated tool retrieval component. Tool retrieval is nontrivial due to the following challenges: 1) complex user instructions and tool descriptions; 2) misalignment between tool retrieval and tool usage models. To address the above issues, we propose to enhance tool retrieval with iterative feedback from the large language model. Specifically, we prompt the tool usage model, i.e., the LLM, to provide feedback for the tool retriever model in multi-round, which could progressively improve the tool retriever’s understanding of instructions and tools and reduce the gap between the two standalone components. We build a unified and comprehensive benchmark to evaluate tool retrieval models. The extensive experiments indicate that our proposed approach achieves advanced performance in both in-domain evaluation and out-of-domain evaluation [^1] Code available at https://github.com/travis-xu/TR-Feedback..
@@
Qiancheng Xu, Yongqi Li, Heming Xia, Wenjie Li Department of Computing, The Hong Kong Polytechnic University, China {qiancheng.xu, he-ming.xia}@connect.polyu.hk liyongqi0@gmail.comcswjli@comp.polyu.edu.hk
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable success in language-related tasks and are considered a potential pathway to achieving artificial general intelligence Zhao et al. (2023). However, despite their powerful capabilities, LLMs are still limited in many aspects, such as knowledge update and mathematical reasoning. A promising way to overcome these limitations is to empower LLMs with external tools, known as tool learning Qin et al. (2023a); Qu et al. (2024a). Tool learning not only enhances LLMs’ performance on existing tasks but also allows them to tackle tasks that were previously beyond their reach. Besides, the ability to use tools is a crucial hallmark on the path to advanced intelligence.
@@
Despite the practicality and necessity, tool retrieval has been inadequately studied. Some approaches have adopted traditional document retrieval methods to retrieve tools for LLMs Li et al. (2023); Qin et al. (2023b). However, we argue that they overlook the unique challenges of tool retrieval for LLMs: 1) Complex usser instructions and tool descriptions. As illustrated in Figure 2, compared with document retrieval, user instructions are usually ambiguous and complex, and the reputation rate between instructions and corresponding tool descriptions is much lower. Unfortunately, the retriever model is typically limited in its capacities because of the efficiency requirements, which makes tool retrieval more difficult and challenging. 2) Misalignment between tool retrieval and tool usage models. Previous approaches deploy the tool retriever separately from the downstream tool-usage model, which hinders the LLM from knowing which tools are really useful from the tool-usage perspective. Thus, it will result in a tool recognition gap between the tool retriever and tool usage model, degrading the tool-use performance further.
@@ 2.1 Tool Learning in LLMs
Tool learning aims to equip LLMs with external tools to enhance and expand their capabilities Ruan et al. (2023); Wang et al. (2024b); Huang et al. (2024c). Generally, existing tool learning methods could be categorized into in-context learning and fine-tuning approaches. The former approach encourages LLMs to use tools with descriptions, documentation, or demonstrations Yuan et al. (2024); Du et al. (2024); Mu et al. (2024), while the latter one trains the parameters of LLMs using specially created tool-use datasets Hao et al. (2023); Tang et al. (2023); Gao et al. (2024). However, no matter whether the in-context learning or fine-tuning approach encounters severe challenges in real-world scenarios, where the candidate tools are extensive and frequently updated. Therefore, it is crucial to equip LLMs with a tool retrieval component to select appropriate tools from a large-scale tool set. Recent works have proposed a stopgap measure through traditional document retrieval Patil et al. (2023); Qin et al. (2023b); Zheng et al. (2024), task decomposition Anantha et al. (2023); Huang et al. (2024b) and graph-based methods Qu et al. (2024b). In this work, we aim to develop a method specialized for enhancing the tool retriever.
@@ 4.1 Overview
Recent studies have found that LLMs show a great capability in acting as a critic Zheng et al. (2023) and could provide comprehensive feedback to improve performance across a range of tasks Madaan et al. (2023); Asai et al. (2023). Inspired by those observations, we propose an innovative framework that leverages the LLM’s feedback to improve the tool retrieval process iteratively. Different from approaches which focus on feedback from execution results after tool execution step Yao et al. (2023); Wang et al. (2024a), we obtain LLMs’ feedback before the actual tool execution step, i.e., right after the tool retrieval step.
@@ 4.2 Feedback Generation
The feedback from the LLM is finalized made on the current user instruction $q^{t}$. Specifically, we prompt the LLM to generate refined instruction with enriched information in two dimensions: 1) more detailed and personalized content about those user’s intent which have not been solved by current tools, helping the retriever explore other relevant tools; (2) more scenario-specific tool-usage information about existing appropriate tools, helping the retriever give higher ranking priority to those tools. This step can be formulated as,
@@ 4.3 Iteration-Aware Feedback Training
We concatenate a special token “Iteration $t$” in front of the instruction, where $t$ is the instruction’s iteration step (e.g., “Iteration $t-1$” for $q^{t-1}$ and “Iteration $t$” for $q^{t}$).
We also employ the hard negative sampling in training. Concretely, for each given instruction, we randomly sample an incorrect tool from the retrieved top-$K$ tool list. The high similarity scores of those tools indicate that they are prone to be mistaken as correct tools by the retriever. In feedback training, we utilize those tool-instruction pairs as hard negative samples. Then the loss function for each iteration could be calculated as,
@@
Then the final training objective could be formulated as the sum of losses in each iteration as follows,
$\mathcal{L}_{feedback}=\sum_{t=1}^{T}\alpha^{t}\mathcal{L}(q^{t}),$ (6)
where $\alpha^{t}$ is a balancing factor and $L(q^{t})$ is the loss function calculated by Equation 5 based on the refined user instructions $q^{t}$ in the $t$th iteration. In this way, the LLM’s comprehensive knowledge of the user requirements could be injected into the retriever through those refined instructions. Besides, with the aid of iteration-aware tokens and joint-training manner, the retriever could maintain a balance between newly learned knowledge and previously acquired knowledge.
@@ 5.1 Setup
Datasets and evaluation. To assess the tool retrieval performance of models, we conduct an experiment on tool retrieval benchmark, referred to as TR-bench, based on three datasets, including ToolBench Qin et al. (2023b), T-Eval Chen et al. (2023), and UltraTools Huang et al. (2024a). To address real-world requirements, we conduct evaluations in both in-domain and out-of-domain settings. Specifically, the training set is from ToolBench, while the test set of ToolBench is employed for in-domain evaluation, and the test sets from T-Eval and UltraTools are used for out-of-domain evaluation. The statistics of TR-bench are summarized in Table 1.
@@
Out-of-domain Evaluation T-Eval 553 50
@@
Table 1: Statistics of the TR-bench, which is conducted from ToolBench Qin et al. (2023b), T-Eval Chen et al. (2023), and UltraTools Huang et al. (2024a).
Following ToolBench, we adopt the Normalized Discounted Cumulative Gain (NDCG) Järvelin and Kekäläinen (2002), an ideal metric for tool retrieval to evaluate the quality of retrieved tools. In our evaluation, we report NDCG@$m$ ($m=1,3,5,10$), calculated according to the position of each golden tool among top-$m$ candidates tools retrieved by the tool retriever. Thus, the more accurately the tool retriever can retrieve correct tools, the higher the NDCG@$m$ score will be.
@@
Table 3: Out-of-domain evaluation on TR-bench in terms of NDCG@$m$ under two scenarios, T-Eval Chen et al. (2023) and UltraTools Huang et al. (2024a). % improve represents the relative improvement achieved by our method over the previously best tool retrieval method.
Implementation details. We employ Sentence-BERT Reimers and Gurevych (2019) to train our retriever model based on BERT-base Devlin et al. (2019). We set the learning rate to $2e{-5}$ with 500 warm-up steps. The batch size in training is set to $64$. We utilize ChatGPT (gpt-3.5-turbo-0125) [^3] https://openai.com/index/introducing-chatgpt-and-whisper-apis/. as the LLM for giving feedback. The number of tool candidates $K$, the balancing factor $\alpha$, and the iteration round $T$ are set to 10, 1, and 3, respectively. We have trained the model several times to confirm that the improvement is not a result of random chance and present the mid one. Our experiments were conducted on four NVIDIA A6000 GPUs with 48 GB of memory.
@@ 5.2 Main Results
We further compare the tool usage performance of our method with ToolRetriever in the I2 scenario. We adopt ToolLLaMA Qin et al. (2023b) which is trained on LLM-annotated solution path as the tool usage model, and use “pass rate” and “win rate” as evaluation metrics. Our method achieves 75.6% for pass rate compared to ToolRetriever’s 68.5%, and 65.9% for win rate compared to ToolRetriever’s 60.8%. The results demonstrates the performance improvement in tool usage, benefiting the entire tool learning process.
@@
Table 5: Analysis on iteration round under the intra-category multi-tool (I2) scenario. The efficiency is measured by the time consumption to complete one user instruction.
@@
Table 6: Analysis on different base models under the intra-category multi-tool (I2) scenario.
@@
Table 7: Analysis on embedding sizes under the intra-category multi-tool (I2) scenario.
5.3 Ablation Study
@@ 6 Conclusion and Future Work
1) Undoubtedly, our iterative refinement will reduce the inference speed of the tool retrieval. The efficiency issue is inherent in approaches involving LLMs’ interaction. We have evaluated the efficiency as the number of iterative rounds increases. Fortunately, we observed that the retrieval model can achieve a significant performance improvement after just a single round of LLMs’ feedback compared to without feedback. Furthermore, the performance enhancement of the tool retrieval is crucial for the subsequent tool usage model, ensuring that the correct tools are retrieved and lays the foundation for all subsequent steps of tool usage. Therefore, we believe that performance improvement is worthwhile despite some efficiency loss. We will also pay more attention to this issue in the future. 2) Similar to document retrieval, the used datasets in our work also contain “false negative” samples. For instance, some tools may be capable of handling the user’s instruction but are not labeled as positive. This can disrupt the training and evaluation of tool retrieval and is a common limitation in many retrieval scenarios.